x²(y-z)p + y²(z − x)q = z²(x-y)


1
Expert's answer
2021-07-28T15:53:51-0400

Given the equation:


"x^2(y-z)p + y^2(z \u2212 x)q = z^2(x-y)"

The equation is a Lagrange Linear PDE and it's of the form:


"Pp+Qq=R"

with:


"P=x^2(y-z); \quad Q=y^2(z − x); \quad R=z^2(x-y)"

The auxiliary equation is in the form:


"\frac{dx}{P}=\frac{dy}{Q}=\frac{dz}{R}"

Thus:


"\frac{dx}{x^2(y-z)}=\frac{dy}{y^2(z − x)}=\frac{dz}{z^2(x-y)}"

We proceed to solve the equation.


Using the multipliers "\frac{1}{x},\frac{1}{y},\frac{1}{z}"


"\frac{\frac{dx}{x}}{x(y-z)}=\frac{\frac{dy}{y}}{y(z − x)}=\frac{\frac{dz}{z}}{z(x-y)}\\"

Adding up we have:


"\frac{dx}{x}+\frac{dy}{y}+\frac{dz}{z}=0"

Integrating through:


"\int\frac{dx}{x}+\int\frac{dy}{y}+\int\frac{dz}{z}=\int0\\
\ln x + \ln y+ \ln z=\ln c\\
\ln(xyz)=\ln c\\
\bold{c_1 = xyz}"

Using the multipliers "\frac{1}{x^2}, \frac{1}{y^2}, \frac{1}{z^2}" as another triplet such that the denominator vanishes:

"\frac{\frac{dx}{x^2}}{y-z}=\frac{\frac{dy}{y^2}}{z − x}=\frac{\frac{dz}{z^2}}{x-y}\\"

Adding up, we get:


"\frac{dx}{x^2}+\frac{dy}{y^2}+\frac{dz}{z^2}=0"

Integrating through:


"\int\frac{dx}{x^2}+\int\frac{dy}{y^2}+\int\frac{dz}{z^2}=\int 0\\
-\frac{1}{x}-\frac{1}{y}-\frac{1}{z}=c_2\\
\bold{c_2=-\Big(\frac{1}{x}+\frac{1}{y}+\frac{1}{z}\Big)}"

Therefore the solution of the PDE is:


"\phi(c_1,c_2) = \Big(xyz, -\Big(\frac{1}{x}+\frac{1}{y}+\frac{1}{z}\Big)\Big)"